OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Numerical Answers Expected!

OpenStudy (anonymous):

it is 22

OpenStudy (anonymous):

Can you explain how? for future reference?

OpenStudy (anonymous):

Directly substitute n = 5 into the formula. f(5) = f(1) + f(2) + f(4) .. Keep this in mind. We know f(1) = 2, and f(2) = 4 so f(5) = 2 + 4 + f(4) = 6 + f(4). Now determine f(4) . . . f(4) = f(1) + f(2) + f(3) = 6 + f(3)... so.. f(5) = 6 + (6 + f(3) ) = 12 + f(3)

OpenStudy (anonymous):

Here is the general rule: f(n) = f(1) + f(2) + f(n - 1)

OpenStudy (anonymous):

We also know the value of f(1) and f(2). f(1) = 2, f(2) = 4. So we can rewrite f(n) as.. f(n) = 2 + 4 + f(n -1) = 6 + f(n - 1).

OpenStudy (anonymous):

You said f(2) = 4 but it says in the problem that f(2) = 3 ?????

OpenStudy (anonymous):

See if you can now determine f(3) knowing that f(n) = 6 + f(n - 1)

OpenStudy (anonymous):

oh, my mistake

OpenStudy (anonymous):

one moment then

OpenStudy (anonymous):

So we have the conditions: f(1)=2f(2)=3f(n)=f(1)+f(2)+f(n−1);n>2 We can simplify f(n) to: f(n)=5+f(n−1) So therefore: f(5)=5+f(5−1)=5+f(4)=5+[5+f(3)]=5+[5+(5+f(2))]=5+5+5+3=18

OpenStudy (anonymous):

You have the following sequence of numbers leading up: f(n)=5+f(n−1);n>2f(1)=2f(2)=3f(3)=8f(4)=13f(5)=18

OpenStudy (anonymous):

its 22

OpenStudy (anonymous):

no, it is 18

OpenStudy (anonymous):

Thanks! And do you mind answering one more problem???

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

suree :3

OpenStudy (anonymous):

f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

Generate the first 5 terms of this sequence:

OpenStudy (anonymous):

I am absolutely lost on this one

OpenStudy (anonymous):

in what order?

OpenStudy (anonymous):

ascending or descending?

OpenStudy (anonymous):

Ascending I assume

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

f(5) = f(4) + f(3) You have to figure out f(3) first. Then figure out f(4) Then figure out f(5)...the 5th term

OpenStudy (anonymous):

Where does it say f(5) ?

OpenStudy (anonymous):

It's the fifth term, find it using that ^ 0, 1, 1, 2, 3

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

each number in the sequence is the sum of the previous two

OpenStudy (anonymous):

medal please?

OpenStudy (anonymous):

I did...

OpenStudy (anonymous):

o otay thanks :3